# Absolute frequency measurement of a Lu+ (3D1) optical frequency standard via link to international atomic time

### v2

@@ Abstract
We report on an absolute frequency measurement of the ${\rm Lu}^{+}\,(^{3}\rm D_{1})$ standard frequency which is defined as the hyperfine-average of ${}^{1}\rm S_{0}$ to ${}^{3}\rm D_{1}$ optical clock transitions in ${}^{176}{\rm Lu}^{+}$. The measurement result of 353 638 794 073 800.34(32)353638794073800.3432353\,638\,794\,073\,800.34(32)353 638 794 073 800.34 ( 32 )Hz with a fractional uncertainty of $9.1\times 10^{-16}$ was obtained by operating a single-ion ${}^{176}{\rm Lu}^{+}$ frequency standard intermittently over 3 months with a total uptime of 162 hours. Traceability to the International System of Units (SI) is realized by remote link to International Atomic Time. This is the first reported absolute frequency value for a ${\rm Lu}^{+}\,(^{3}\rm D_{1})$ optical frequency standard.
@@ 2.1 ${}^{176}{\rm Lu}^{+}$ optical frequency standard
The choice of Ramsey time and number of cycle between servo updates is limited by the clock laser stability. For this work the Ramsey time was chosen conservatively to ensure robust operation in which the servo never slips to an incorrect Ramsey fringe. The quantum projection noise limited fractional frequency instability during clock operation is
$\sigma(\tau)=\frac{1}{f_{\mathrm{Lu}}}\frac{1}{2\pi CT_{R}\sqrt{2N}}\sqrt{\frac{t_{u}}{\tau}}\approx\frac{3.9\times 10^{-15}}{\sqrt{\tau}},$
@@ 2.2 Link to SI second
Table 1: Nominal uncertainty budget of Lu clock for operating parameters used in this campaign.
Effect Correction Uncertainty
($10^{-18}$) ($10^{-18}$)
Quadradic Zeeman shift -140.7 0.6
Excess Micromotion 0.2 0.2
Second-order Doppler (thermal) -0.1 0.1
@@
Total -150.0 7.3
Gravitational redshift 1718. 36.
Total (with redshift) 1568. 37.
@@ 3 Lu+ Systematics
The Lu+ clock systematic uncertainty budget for the operating parameters of this measurement campaign is summarised in table 1. Data presented in this work was collected using the apparatus ‘Lu-2’ shortly after the clock comparison and systematics evaluation reported in [8]. Compared to the uncertainty budget in [8], probe induced systematic shifts are larger here due to teh shorter Ramsey interrogation time. In [8] two atomic references were compared using correlation spectroscopy which enabled interrogation times beyond the coherence time of the laser. The only frequency corrections $>10^{-17}$ and of statistical significance for the absolute frequency assessment are from the quadratic Zeeman shift and gravitational redshift. Remaining systematics are only briefly discussed.
@@ 3.3 Other Systematics
Excess micromotion was measured in three orthogonal directions on three occasions throughout the measurement campaign using sideband spectroscopy [28] on the 804 nm $|\,^{1}S_{0},7,0\,\rangle$ to $|\,^{3}D_{2},9,0\,\rangle$ clock transition. In each instance the corresponding second-order Doppler shift was $<2\times 10^{-19}$ before (re-)optimizing the compensation. it is noted that an improved micromotion compensation method was developed after this work [11].
The Second-order Doppler shift due to thermal motion and the ac-Zeeman shift due to rf fields at the trap drive frequency are as evaluated in [8]. The microwave ac-Zeeman shift arises from imbalanced $\sigma^{\pm}$ polarisation components of the microwave fields which off-resonantly couple to $\Delta m=\pm 1$ Zeeman transitions. The polarisation of both microwave fields was evaluated three times throughout the campaign by measuring the relative coupling strengths of $\Delta m=[-1,0,1]$ microwave transitions. The observed coupling strengths did not vary by more the 1% indicating stable microwave polarizations throughout. One of the microwave fields had imbalanced $\sigma^{\pm}$ components resulting in the uncharacteristically large correction and uncertainty due to this effect.
An error in microwave coupling strengths relative to the values determined by their respective $\pi$ time gives rise to an imperfect HA [25]. The uncertainty in table 1 corresponds to a 1% error in the microwave coupling strengths.
For hyper-Ramsey spectroscopy [15] the clock laser is frequency stepped by the evaluated ac-Stark shift during the clock interrogation pulses. The remaining clock shift after hyper-Ramsey, in Hz, is given by $\frac{2}{\pi T_{R}}\left(\frac{\Delta}{\Omega}\right)^{3}$ where $\Delta$ is the error in the frequency step and $\Omega=\pi/\tau_{L}$ is the optical coupling strength. The ac-Stark shift is evaluated to $\sim$1precision prior to clock operation by interleaved comparison of hyper-Ramsey and Rabi spectroscopy. Accounting for possible variation in laser intensity over the course of the clock operation, we allow for 5% error in the ac-Stark shift step in estimating the uncertainty for table 1.
@@
We note that the systematic uncertainty reported here is evaluated for the operating parameter of this campaign. These parameters were not optimised to reduce systematic uncertainty beyond the requirements of the absolute frequency measurement. The microwave ac-Zeeman shift can be suppressed by balancing the $\sigma^{\pm}$ polarisation components, or using a longer microwave $\pi$ time. The ac-Stark shift can be suppressed by using a longer optical pulse duration. With improved laser stability, the Ramsey time can be extended to reduce all probe induced systematics effects, which includes the microwave coupling, the microwave ac-Zeeman, the clock laser ac-Stark systematics.
@@ 4.1 Maser drift
The hydrogen maser drift is evaluated relative to UTC(USNO) over an 80 day interval encompassing the optical frequency measurements, as shown in figure 4. Over this interval the maser frequency drift slightly deviates from linear so a quadratic model is used. The maser linear drift over this is interval is near $-3.7\times 10^{-15}\,\mathrm{d}^{-1}$ and aging at a rate $4.6\times 10^{-15}\,\mathrm{d}^{-1}\mathrm{year}^{-1}$. The fitted maser drift model is subtracted from the local observations, the blue points in figure 4(a), to link the measurements to UTC(USNO) up to uncertainty from the additional noise process on the maser frequency which are assumed to have zero mean. The blue points in figure 4(b), which are $-y(\mathrm{Lu}/\mathrm{UTC(USNO)})$, are binned in 5 day intervals aligned to Circular T and averaged. The uncertainty in the 5 d averages of $y(\mathrm{Lu}/\mathrm{UTC(USNO)})$ has three contributions. First, the statistical uncertainty in the five day averages of $f_{\mathrm{HM}}/f_{\mathrm{UTC(USNO)}}$ relative to the maser drift model fit on the the 80 day interval. This is estimated to be $1.5\times 10^{-15}$ from the instability of the residuals, the orange points in figure 4(b), at five days, and is consistent with the Hadamard deviation of $f_{\mathrm{HM}}/f_{\mathrm{UTC(USNO)}}$ which is shown in figure 3. Second, the fitted drift model contributes uncertainty which is correlated across all measurements, and is estimated to contribute $2\times 10^{-16}$ to the final result. Third, the uncertainty from measurement dead time which is evaluated in the following section.
4.2 Dead time uncertainty
@@
The dead time uncertainty is calculated by simulation of the maser noise [31, 32]. A maser noise model is constructed to be consistent with both the observed $f_{\mathrm{Lu}}/f_{\mathrm{HM}}$ instability on shorter timescales, and the maser flicker noise floor inferred from GPS link to UTC(USNO) on longer times scales. The allan deviation for the longest continuous measurement of $f_{\mathrm{Lu}}/f_{\mathrm{HM}}$ (12 h) is shown in figure 3. The noise model is the sum of white phase noise $5.1\times 10^{-13}\,(\tau/\mathrm{s})^{-1}$, white frequency noise $4.6\times 10^{-14}\,(\tau/\mathrm{s})^{-1/2}$, flicker frequency noise $1.3\times 10^{-15}$, and the bandwidth limited noise labeled ‘plateau’ in figure 3 which reproduces the instability plateau observed for averaging times $\lesssim 1\,\mathrm{h}$. The plateau has however been consistently observed in $f_{\mathrm{Lu}}/f_{\mathrm{HM}}$ measurements over the last three years. The generated noise [33] is sampled to simulate $\sim 2\times 10^{3}$ trajectories on the interval $T_{2}=5\,\mathrm{d}$. For an uptime $T_{1}$, which may be distributed non-continuously over the interval $T_{2}$, the dead time uncertainty is calculated from the standard deviation of the difference in averages over $T_{1}$ and $T_{2}$. For the actual uptimes $T_{1}$ in the experimental data, the evaluated dead time uncertainties are shown as the black points in figure 5. The influence of the distribution of $T_{1}$ was further investigated by sampling random arrangements of $T_{1}$ as a function of total uptime fraction, with the restriction of a minimum (maximum) continuous experiment run time of 1 (12) hours to mimic typical operation. The gray shaded region in figure 5 bounds the maximum and minimum dead time uncertainty found under these criteria. The solid lines correspond to two special cases, $T_{1}$ as a single continuous run centered on $T_{2}$ (blue) and $T_{1}$ divided into five runs evenly spaced over the five days (orange). For future measurement campaigns this simulation can serve as a practical guide to determine the uptime fraction and dead-time distribution required to achieve a target uncertainty.
@@ 4.3 Diurnal effects
Correlations with a daily temperature cycle are of particular concern because the majority of the optical clock uptime was during the daytime. We analyzed all measurement residuals after maser drift subtraction for correlation with time of day acquired, but there is no definitive diurnal amplitude. Statistically, we can only bound the amplitude to $<1\times 10^{-15}$ and at this upper bound the final result would be biased by $5\times 10^{-16}$ when accounting for the distribution of uptime. We include this potential bias as a systematic uncertainty in the final result.
Figure 5: Estimated dead time uncertainty from simulation. Black points are simulated dead time uncertainty for the uptimes distributions present in the experimental data. See main text for description of other simulated results shown.
@@ 4.4 Link to TAI
Ratio Uncertainty ×$\times 10^{-16}$
@@ 4.5 Link to SI
Table 3: Uncertainty budget for the absolute frequency measurement of the ${}^{176}\mathrm{Lu}^{+}$ ${}^{1}\rm S_{0}-{{}^{3}}\rm D_{1}$ frequency standard for each month, and the total campaign. All uncertainties given in $10^{-16}$ fractional frequency uncertainty.
@@
Lu/SI Total Unc. 11 14 20 09.1
Figure 6: The measured frequency offset for each month (black points) relative to the weighted mean value. Black error bars are the total uncertainties, and red error bars the statistical component only. The weighted mean of all three months is 353 638 794 073 800.34(32)353638794073800.3432353\,638\,794\,073\,800.34(32)\,353 638 794 073 800.34 ( 32 )Hz, with the fractional uncertainty of $9.1\times 10^{-16}$ indicated by the gray shaded region.
@@ 5 Discussion
The absolute frequencies obtained for each circular T report are shown in figure 6. The result of 353 638 794 073 800.34(32)353638794073800.3432353\,638\,794\,073\,800.34(32)\,353 638 794 073 800.34 ( 32 )Hz is obtained from the weighted mean of values from the three months. The reduced chi squared is $\chi^{2}_{\nu}=0.56$ with $\nu=2$ degrees of freedom.
The fractional uncertainty of $9.1\times 10^{-16}$ is limited by the satellite link stability and dead time uncertainty. For future campaigns, improved satellite link stability might be achieved by PPP with carrier-phase integer ambiguity resolution (PPP-AR) [37, 38]. Processing of more recent GNSS data with the PPP-AR algorithm implemented by the Natural Resources Canada (NRCan) online service shows an instability of $\sim 1\times 10^{-15}$ at 1 d averaging time, compared to the 5 days observed in this work. To realize this improvement it was also necessary to relocate the rooftop GNSS aerial away from a nearby building partially obstructing the sky.
The dead time uncertainty may be reduced by either higher uptime of the optical clock or improved stability of the local frequency reference. During preparation of this report, 95% clock uptime was achieved over ten consecutive days. The primary source of clock interruptions remaining is mode hopping of a 701 nm laser which is frequency doubled to generate the 350 nm repump light. An alternative solution for generating 350 nm light by third harmonic generation is in progress. If this proves a robust solution, uptime approaching 100% appears to be imminently feasible, limited only by interruption due to ion loss. Our trapping lifetime for a single ion is typically several days in both experiment chambers.
A stabilized optical fiber link to the Singapore National Metrology Centre (NMC) has been recently established and will enable linking of our hydrogen masers. One the hydrogen masers of NMC is specified to have much better stability than the one used here, but this is yet to be verified against our optical clock. An ensemble of hydrogen masers could provide an improved stability flywheel and better characterisation of the maser noise of individual maser by continuous observation of the ensemble [39].
@@
The authors thank Nils Nemitz for critical reading of the manuscript and Alex Voss for orthometric levelling measurements to determine the ellipsoidal height. The RINEX files of USN7 were obtained through the online archives of the Crustal Dynamics Data Information System (CDDIS), NASA Goddard Space Flight Center, Greenbelt, MD, USA. https://cddis.nasa.gov/archive/gnss/data/daily/ This research is supported by the National Research Foundation (NRF), Singapore, under its Quantum Engineering Programme (QEP-P5); the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme (NRF2021-QEP2-01-P03) and through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009); and the Ministry of Education, Singapore under its Academic Research Fund Tier 2 (MOE-T2EP50120-0014).

### v3

@@ Abstract
We report on an absolute frequency measurement of the ${\rm Lu}^{+}\,(^{3}\rm D_{1})$ standard frequency which is defined as the hyperfine-average of ${}^{1}\rm S_{0}$ to ${}^{3}\rm D_{1}$ optical clock transitions in ${}^{176}{\rm Lu}^{+}$. The measurement result of 353 638 794 073 800.35(33)353638794073800.3533353\,638\,794\,073\,800.35(33)353 638 794 073 800.35 ( 33 )Hz with a fractional uncertainty of $9.2\times 10^{-16}$ was obtained by operating a single-ion ${}^{176}{\rm Lu}^{+}$ frequency standard intermittently over 3 months with a total uptime of 162 hours. Traceability to the International System of Units (SI) is realized by remote link to International Atomic Time. This is the first reported absolute frequency value for a ${\rm Lu}^{+}\,(^{3}\rm D_{1})$ optical frequency standard.
@@ 2.1 ${}^{176}{\rm Lu}^{+}$ optical frequency standard
The choice of Ramsey time and number of cycles between servo updates is limited by the clock laser stability. For this work the Ramsey time was chosen conservatively to ensure robust operation in which the servo never slips to an incorrect Ramsey fringe. The quantum projection noise limited fractional frequency instability during clock operation is
$\sigma(\tau)=\frac{1}{f_{\mathrm{Lu}}}\frac{1}{2\pi CT_{R}\sqrt{2N}}\sqrt{\frac{t_{u}}{\tau}}\approx\frac{3.9\times 10^{-15}}{\sqrt{\tau}},$ (2)
@@ 2.2 Link to SI second
Table 1: Nominal uncertainty budget of the Lu+ clock for operating parameters used in this campaign. All shifts and uncertainties are given fractionally relative to the 353 THz ${\rm Lu}^{+}\,(^{3}\rm D_{1})$ standard frequency.
Shift Uncertainty
Effect (×$\times 10^{-18}\,\mathrm{Hz/Hz}$) (×$\times 10^{-18}\,\mathrm{Hz/Hz}$)
Quadratic Zeeman shift -140.7 0.6
Excess micromotion -0.2 0.2
Second-order Doppler -0.1 0.1
@@
Total -150.4 7.3
Gravitational redshift 1718 36
Total (with redshift) 1568 37
@@ 3 Lu+ Systematics
The Lu+ clock systematic uncertainty budget for the operating parameters of this measurement campaign is summarised in table 1. Data presented in this work was collected using the apparatus ‘Lu-2’ shortly after the clock comparison and systematics evaluation reported in [8]. Compared to the uncertainty budget in [8], probe induced systematic shifts are larger here due to the shorter Ramsey interrogation time. In [8] two atomic references were compared using correlation spectroscopy which enabled interrogation times beyond the coherence time of the laser. The only frequency corrections $>10^{-17}$ and of statistical significance for the absolute frequency assessment are from the quadratic Zeeman shift and gravitational redshift. Remaining systematics are only briefly discussed.
@@ 3.3 Other Systematics
Excess micromotion was measured in three orthogonal directions on three occasions throughout the measurement campaign using sideband spectroscopy [28] on the 804 nm $|\,^{1}S_{0},7,0\,\rangle$ to $|\,^{3}D_{2},9,0\,\rangle$ clock transition. In each instance the magnitude of the corresponding second-order Doppler shift was $<2\times 10^{-19}$ before (re-)optimizing the compensation. It is noted that an improved micromotion compensation method was developed after this work [11].
The Second-order Doppler shift due to thermal motion and the ac-Zeeman shift due to rf fields at the trap drive frequency are as evaluated in [8]. The microwave ac-Zeeman shift arises from imbalanced $\sigma^{\pm}$ polarisation components of the microwave fields which off-resonantly couple to $\Delta m=\pm 1$ Zeeman transitions. The polarisations of both microwave fields were evaluated three times throughout the campaign by measuring the relative coupling strengths of $\Delta m=[-1,0,1]$ microwave transitions. The observed coupling strengths did not vary by more the 1% indicating stable microwave polarizations throughout. One of the microwave fields had imbalanced $\sigma^{\pm}$ components resulting in an uncharacteristically large shift and uncertainty, $-8.0(6.2)\times 10^{-18}$, due to this effect.
An error in microwave coupling strengths relative to the values determined by their respective $\pi$ time gives rise to imperfect hyperfine averaging [25]. The uncertainty in table 1 corresponds to a 1% error in the microwave couplings.
For hyper-Ramsey spectroscopy [15] the clock laser is frequency stepped by the evaluated ac-Stark shift during the clock interrogation pulses. The remaining clock shift after hyper-Ramsey, in Hz, is given by $\frac{2}{\pi T_{R}}\left(\frac{\Delta}{\Omega}\right)^{3}$ where $\Delta$ is the error in the frequency step and $\Omega=\pi/\tau_{L}$ is the optical coupling strength. The ac-Stark shift is evaluated to $\sim$1precision prior to clock operation by interleaved comparison of hyper-Ramsey and Rabi spectroscopy. Accounting for possible variation in laser intensity over the course of clock operation, we allow for 5% error in the ac-Stark shift step in estimating the uncertainty for table 1.
@@
We note that the systematic uncertainty reported here is evaluated for the operating parameter of this campaign. These parameters were not optimised to reduce systematic uncertainty beyond the requirements of the absolute frequency measurement. The microwave ac-Zeeman shift can be suppressed by balancing the $\sigma^{\pm}$ polarisation components, or using a longer microwave $\pi$ time. The ac-Stark shift can be suppressed by using a longer optical pulse duration. With improved laser stability, the Ramsey time can be extended to reduce all probe induced systematics effects, which includes the microwave coupling, the microwave ac-Zeeman, and the (848 nm) ac-Stark effects.
@@ 4.1 Maser drift
The hydrogen maser drift is evaluated relative to UTC(USNO) over an 80 day interval encompassing the optical frequency measurements, as shown in figure 4. Over this interval the maser frequency drift slightly deviates from linear so a quadratic model is used. The maser linear drift over this is interval is near $-3.7\times 10^{-15}\,\mathrm{d}^{-1}$ and aging at a rate $4.6\times 10^{-15}\,\mathrm{d}^{-1}\mathrm{year}^{-1}$. The fitted maser drift model is subtracted from the local observations, the blue points in figure 4(a), to link the measurements to UTC(USNO) up to uncertainty from the additional noise process on the maser frequency which are assumed to have zero mean. The blue points in figure 4(b), which are $-y(\mathrm{Lu}/\mathrm{UTC(USNO)})$, are binned in 5 day intervals aligned to Circular T and averaged. The uncertainty in the 5 d averages of $y(\mathrm{Lu}/\mathrm{UTC(USNO)})$ has three contributions. First, the statistical uncertainty in the five day averages of $f_{\mathrm{HM}}/f_{\mathrm{UTC(USNO)}}$ relative to the maser drift model fit on the 80 day interval. This is estimated to be $1.5\times 10^{-15}$ from the instability of the residuals, the orange points in figure 4(b), at five days, and is consistent with the Hadamard deviation of $f_{\mathrm{HM}}/f_{\mathrm{UTC(USNO)}}$ which is shown in figure 3. Second, the fitted drift model contributes uncertainty which is correlated across all measurements, and this is estimated to contribute $2\times 10^{-16}$ to the final result. Third, the uncertainty contributed from interpolation of the HM over optical clock dead time, which is evaluated in the following section.
4.2 HM interpolation uncertainty
The uncertainty due to HM interpolation over optical clock dead time is estimated by simulation of the maser noise [31, 32]. A maser noise model is constructed to be consistent with both the observed $f_{\mathrm{Lu}}/f_{\mathrm{HM}}$ instability on timescales less than a few hours, and the maser flicker noise floor inferred from GPS link to UTC(USNO) on timescales longer than five days. The Hadamard deviation for the longest continuous measurement of $f_{\mathrm{Lu}}/f_{\mathrm{HM}}$ (12 h) is shown in figure 3. The noise model is the sum of white phase noise $5.1\times 10^{-13}\,(\tau/\mathrm{s})^{-1}$, white frequency noise $4.6\times 10^{-14}\,(\tau/\mathrm{s})^{-1/2}$, flicker frequency noise $1.3\times 10^{-15}$, and the additional noise labeled ‘plateau’ in figure 3 which reproduces the instability plateau observed for averaging times in the interval $\sim$5-60 minutes. This plateau clearly cannot be modelled by only power law noise but is captured reasonably here with the addition of low-pass filtered white frequency noise. The plateau has been consistently observed in $f_{\mathrm{Lu}}/f_{\mathrm{HM}}$ measurements over the last three years.
@@ 4.2 HM interpolation uncertainty
Power law noise is generated by fast Fourier transform (FFT) methods [33]. For the ‘plateau’ noise, a high frequency cutoff is applied before the FFT to achieve low pass filtering. The total noise is generated by summing the time series generated for the individual processes. The total noise is sampled to simulate $\sim 2\times 10^{3}$ trajectories on the interval $T_{2}=5\,\mathrm{d}$. For an uptime $T_{1}$, which may be distributed non-continuously over the interval $T_{2}$, the dead time uncertainty is calculated from the standard deviation of the difference in averages over $T_{1}$ and $T_{2}$. For the actual uptimes $T_{1}$ in the experimental data, the evaluated dead time uncertainties are shown as the black points in figure 5. The influence of the distribution of $T_{1}$ was further investigated by sampling random arrangements of $T_{1}$ as a function of total uptime fraction, with the restriction of a minimum (maximum) continuous experiment run time of 1 (12) hours to mimic typical operation. The gray shaded region in figure 5 bounds the maximum and minimum dead time uncertainty found under these criteria. The solid lines correspond to two special cases, $T_{1}$ as a single continuous run centered on $T_{2}$ (blue) and $T_{1}$ divided into five runs evenly spaced over the five days (orange). For future measurement campaigns this simulation can serve as a practical guide to determine the uptime fraction and distribution of dead time required to achieve a target uncertainty.
@@ 4.3 Diurnal effects
Correlations with a daily temperature cycle are of particular concern because the majority of the optical clock uptime was during the daytime. We analyzed all measurement residuals after maser drift subtraction for correlation with the time of day, but no definitive diurnal amplitude was observed. Statistically, we can only bound the amplitude to $<1\times 10^{-15}$ and at this upper bound the final result would be biased by $5\times 10^{-16}$ when accounting for the distribution of uptime. We include this potential bias as a systematic uncertainty in the final result.
Figure 5: Result of simulation to estimate the uncertainty due to HM interpolation over optical clock dead time. Black points are uncertainties estimated by simulating the uptimes distributions present in the experimental data. See main text for description of other simulated results shown.
@@ 4.4 Link to TAI
Ratio Uncertainty (×$(\times 10^{-16}\,\mathrm{Hz/Hz})$
@@ 4.5 Link to SI
Table 3: Uncertainty budget for the absolute frequency measurement of the ${}^{176}\mathrm{Lu}^{+}$ ${}^{1}\rm S_{0}-{{}^{3}}\rm D_{1}$ frequency standard for each month, and the total campaign.
@@
Lu/SI Total Unc. 11 14 20 09.2
Figure 6: The measured frequency offset for each month (black points) relative to the weighted mean value. Black error bars are the total uncertainties, and red error bars the statistical component only. The weighted mean of all three months is 353 638 794 073 800.35(33)353638794073800.3533353\,638\,794\,073\,800.35(33)\,353 638 794 073 800.35 ( 33 )Hz, with the fractional uncertainty of $9.2\times 10^{-16}$ indicated by the gray shaded region.
@@ 5 Discussion
The absolute frequencies obtained for each circular T report are shown in figure 6. The result of 353 638 794 073 800.35(33)353638794073800.3533353\,638\,794\,073\,800.35(33)\,353 638 794 073 800.35 ( 33 )Hz is obtained from the weighted mean of values from the three months. The reduced chi squared is $\chi^{2}_{\nu}=0.56$ with $\nu=2$ degrees of freedom.
The fractional uncertainty of $9.2\times 10^{-16}$ is limited by the satellite link stability and HM interpolation uncertainty. For future campaigns, improved satellite link stability might be achieved by PPP with carrier-phase integer ambiguity resolution (PPP-AR) [37, 38]. Processing of more recent GNSS data with the PPP-AR algorithm implemented by the Natural Resources Canada (NRCan) online service shows an instability of $\sim 1\times 10^{-15}$ at 1 d averaging time, compared to the 5 days observed in this work. To realize this improvement it was also necessary to relocate the rooftop GNSS aerial away from a nearby building partially obstructing the sky.
The HM interpolation uncertainty may be reduced by either higher uptime of the optical clock or improved stability of the local frequency reference. During preparation of this report, 95% clock uptime was achieved over ten consecutive days. The primary source of clock interruptions remaining is mode hopping of a 701 nm laser which is frequency doubled to generate the 350 nm repump light. An alternative solution for generating 350 nm light by third harmonic generation is in progress. If this proves a robust solution, uptime approaching 100% appears to be imminently feasible, limited only by interruption due to ion loss. Our trapping lifetime for a single ion is typically several days in both experiment chambers.
A stabilized optical fiber link to the Singapore National Metrology Centre (NMC) has been recently established and will enable linking of our hydrogen masers. One of the hydrogen masers of NMC is specified to have better stability than the one used here, but this is yet to be verified against our optical clock. An ensemble of hydrogen masers could provide an improved stability flywheel and better characterisation of the noise of the individual masers by continuous observation relative to the ensemble [39].
@@
The authors acknowledge Zhiqiang Zhang for contributions to initial data collection and curation. The authors thank Nils Nemitz for critical reading of the manuscript and Alex Voss for orthometric levelling measurements to determine the ellipsoidal height. This research is supported by the National Research Foundation (NRF), Singapore, under its Quantum Engineering Programme (QEP-P5); the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme (NRF2021-QEP2-01-P03) and through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009); and the Ministry of Education, Singapore under its Academic Research Fund Tier 2 (MOE-T2EP50120-0014).
Data availability
The RINEX files of USN7 were obtained through the online archives of the Crustal Dynamics Data Information System (CDDIS), NASA Goddard Space Flight Center, Greenbelt, MD, USA. https://cddis.nasa.gov/archive/gnss/data/daily/.
Data sets that support the findings of this study are available upon reasonable request from the corresponding author.